A parallel-plate capacitor, having plates of 100c㎡ and has three

dielectrics 1mm each and of permittivity’s 3, 4 and 6. If the peak

voltage 2000V is applied to the plates, calculate the energy stored

in each dielectric


1
Expert's answer
2021-10-07T03:15:13-0400

"V=100\cdot10^{-4}\cdot1\cdot10^{-3}=10^{-5}\ (m^3)"


"E=2000/0.003=666667\ (V/m)"


"E_1=\epsilon_1\epsilon_0E^2V/2=3\cdot8.85\cdot10^{-12}\cdot666667^2\cdot 10^{-5}/2="


"=5.9\cdot10^{-5}\ (J)"


"E_2=\epsilon_2\epsilon_0E^2V/2=4\cdot8.85\cdot10^{-12}\cdot666667^2\cdot 10^{-5}/2="


"=7.84\cdot10^{-5}\ (J)"


"E_3=\epsilon_3\epsilon_0E^2V/2=6\cdot8.85\cdot10^{-12}\cdot666667^2\cdot 10^{-5}/2="


"=11.8\cdot10^{-5}\ (J)"